OpenStudy (anonymous):

Projectile Motion Problem: An item is launched from a height of 12 m and its goal is to hit another item 2000 m away and at a height of 3 m. Its velocity is 1000 m/s. What angles could be used to hit it?

OpenStudy (anonymous):

This is kind of what I've been doing so far: \[x _{f} = x _{o} + V _{o}*\cos \alpha *t _{f}\] \[2000 = 1000\cos \alpha * t _{f}\]

OpenStudy (anonymous):

\[y _{f} = y _{o} + V _{o}\sin \alpha *t _{f} + .5*a*t _{f}^2\] \[0 = 9 + 1000\sin \alpha*t _{f} - 4.9t _{f}^2\]

OpenStudy (anonymous):

-h=usin(theta)t-1/2gt^2